# On vector-valued functional equations with multiple recursive terms

### v2

@@ Abstract
In this work, we study vector-valued functional equations with multiple recursive terms that arise naturally when we are dealing with vector-valued multiplicative Lindley-type recursions. Our work is strongly motivated by a wide range of semi-Markovian queueing, and vector-valued autoregressive processes. In all cases, we provide explicit expressions for the joint distribution in terms of Laplace-Stieltjes transforms/generating functions.
Keywords: Recursion; Laplace-Stieltjes transform; Generating functions; Markov-modulated arrivals and services; System of Wiener-Hopf type equations
@@ 1 Introduction
The primary aim of this work is to investigate vector-valued functional equations of the form
@@
for $Re(s)\geq 0$, $Re(\eta)\geq 0$, $|r|<1$, where $\alpha_{i}(s)$, $i=1,\ldots,N$ are commutative contraction mappings, and $\tilde{Z}(r,s,\eta):=(Z_{1}(r,s,\eta),Z_{2}(r,s,\eta),\ldots,Z_{N}(r,s,\eta))^{T}$. Moreover, $G(r,s,\eta)$, $K(r,s,\eta)$ are given vector-valued functions and $\tilde{P}^{(i)}:=(\tilde{P}^{(i)})_{p,q}$, $i,p,q\in E$ is an $N\times N$ matrix, with the element $\tilde{P}^{(i)}_{i,i}=1$, and all the other elements $\tilde{P}^{(i)}_{p,q}=0$, $p,q\neq i$. Note that $\sum_{i=1}^{N}\tilde{P}^{(i)}=I$, that is the identity matrix. Such type of functional equations arise naturally when we are dealing with the time-dependent behaviour of vector-valued multiplicative Lindley-type recursions of a certain type. The corresponding stationary version results in the following vector-valued functional equation:
where $H(s)$, $V(s)$, are known matrix/vector-valued functions. A more general version of (2) is considered in Section 8. We also attempt to consider multidimensional versions such as
@@
which arise in Section 7, and corresponds to the joint LST of a Markov-modulated tandem network of two queue with Levy input and consumption, where $K(s,t)$, $R(s,t)$, $T(s)$, are known matrix-valued functions. Moreover, we also focus on vector-valued reflected autoregressive processes that are described by recursions of the form $\tilde{Z}_{n+1}=[A\tilde{Z}_{n}+\tilde{S}_{n}-\tilde{A}_{n}]^{+}$, where now $A$ is an $N\times N$ known matrix with elements $a_{i,j}\in[0,1)$, and our main concern is to provide expressions for the joint LST of the stationary vector $\tilde{Z}$, as a solution of a functional equation of the form (for $N=2$):
$\begin{array}[]{rl}f(s_{1},s_{2})=&c_{1}(s_{1},s_{2})f(\sum_{j=1}^{2}a_{j,1}s_{j},\sum_{j=1}^{2}a_{j,2}s_{j})-c_{2}(s_{1},s_{2})f(s_{1}a_{1,1}+\lambda_{2}a_{2,1},s_{1}a_{1,2}+\lambda_{2}a_{2,2})\vskip 5.69054pt\\ &-c_{3}(s_{1},s_{2})f(\lambda_{1}a_{1,1}+s_{2}a_{2,1},\lambda_{1}a_{1,2}+s_{2}a_{2,2})+c_{4}(s_{1},s_{2})f(\sum_{j=1}^{2}a_{j,1}\lambda_{j},\sum_{j=1}^{2}a_{j,2}\lambda_{j}),\end{array}$ (4)
where $c_{i}(s_{1},s_{2})$, $i=1,2,3,4$ are known functions.
Our work is strongly motivated by a wide range of semi-Markovian queueing, and reflected autoregressive processes. In particular, (1) corresponds to a vector-valued analogue of equation (2) in [2]. Our primary aim in this work is to consider matrix generalizations of the seminal works in [2, 12]. Quite recently, the author in [20] studied vector-valued functional equations of the form
$\begin{array}[]{rl}\tilde{Z}(r,s,\eta)=&G(r,s,\eta)\tilde{Z}(r,\alpha(s),\eta)+K(r,s,\eta),\vskip 5.69054pt\\ \tilde{Z}(s)=&H(s)\tilde{Z}(a(s))+\tilde{V}(s),\end{array}$ (5)
where $\alpha(s)=as$, $a\in(0,1)$.
In this work, our aim is to extend the analysis in [20] to the case where the focus is on vector-valued functional equations of the form given in (5). Thus, we also extend the analysis in [2] to the matrix-valued framework. Functional equations of the form in (1), (5) arise naturally from the analysis of vector-valued recursions of the form
$\tilde{Z}_{n+1}=[R_{n}(X_{n})\tilde{Z}_{n}+\tilde{Y}_{n}(X_{n})-\tilde{B}_{n}(X_{n})]^{+},$ (6)
where $\{X_{n};n\in\mathbb{N}\}$ denotes an irreducible Markov chain with a finite state space $\{1,2,\ldots,N\}$, and vectors $\tilde{Z}_{n},\tilde{Y}_{n}(X_{n}),\tilde{B}_{n}(X_{n})\in\mathbb{R}^{n}$, and where $R_{n}(X_{n}),\tilde{Y}_{n}(X_{n}),\tilde{B}_{n}(X_{n})$ depend on the state of $X_{n}$, i.e., we consider Markov-modulated vector-valued recursions.
@@ 1.1 Related work
In the following, we provide a brief overview of the existed analytical results in the scalar case, since to our best knowledge, the only vector-valued version of (6) that has been treated analytically refers to the case where $R_{n}(X_{n})=a\in(0,1)$; see [20].
@@
In [12], the authors considered the (scalar) recursion $Z_{n+1}=[aZ_{n}+S_{n}-A_{n+1}]^{+}$ (with $x]^{+}:=max\{x,0\}$), where $\{S_{n}-A_{n+1}\}_{n\in\mathbb{N}_{0}}$ forms a sequence of independent and identically distributed (i.i.d.) random variables and $a\in(0,1)$. The authors provided explicit results for the case where $\{S_{n}\}_{n\in\mathbb{N}_{0}}$ being a sequence of independent $exp(\lambda)$ distributed random variables, and $\{Y_{n}\}_{n\in\mathbb{N}_{0}}$ i.i.d., nonnegative and independent of $\{B_{n}\}_{n\in\mathbb{N}_{0}}$ with distribution function $F_{Y}(.)$ and Laplace–Stieltjes transform (LST) $\phi_{Y}(.)$. Note that in such a case, $Z_{n}$ could be interpreted as the workload in a queueing system just before the $n$th arrival, which adds $Y_{n}$ work, and makes obsolete a fixed fraction 1a1-a of the already present work. The case where $a=1$ corresponds to the classical Lindley recursion describing the waiting time of the classical G/G/1 queue, while the case where a=1a=-1 was investigated in [32]. Further progress has recently been made in [10] where the scalar autoregressive process described by the recursion $Z_{n+1}=[V_{n}Z_{n}+S_{n}-A_{n+1}]^{+}$ was investigated. In [11], the authors considered the case where $V_{n}W_{n}$ was replaced by $F(W_{n})$, where $\{F(t)\}$ is a Levy subordinator (recovering also the case in [12], where $F(t)=at$). Recently, in [8], the authors motivated by applications that arise in queueing and insurance risk models, they considered Lindley-type recursions where the sequences $\{S_{n}\}_{n\in\mathbb{N}_{0}}$, $\{A_{n}\}_{n\in\mathbb{N}_{0}}$ obey a semi-linear dependence. These recursions can also be treated as of autoregressive type. Furthermore, the authors in [2] developed a method to study functional equations that arise in a wide range of queueing, autoregressive and branching processes. Finally, the author in [22], considered a generalized version of the model in [10], by assuming $V_{n}$ to take values in $(-\infty,1]$. In [3, eq. (9)], the authors considered a queueing system with two classes of impatient customers, leading to a specific example of (1) that was analyzed in detail; see also [24]. Quite recently, in [19] the authors generalized the work in [12] by considering, among others, non-trivial dependence structures among $\{S_{n}\}_{n\in\mathbb{N}_{0}}$, $\{A_{n}\}_{n\in\mathbb{N}}$. We also mention the work in [9] where, among others, the authors attempted to provide analytical results for ASIP (asymmetric inclusion process) tandem queues with consumption, the analysis of which leads to a functional equation that is similar to those considered in reflected autoregressive processes.
@@
A primary motivation for our work is related to recent results on vector-valued reflected autoregressive processes. Quite recently, in [20], the author investigated vector-valued reflected autoregressive processes, where the sequences $\{S_{n}\}_{n\in\mathbb{N}_{0}}$, $\{A_{n}\}_{n\in\mathbb{N}}$ are governed by an irreducible background Markovian process with finite state space, i.e., he considered Markov-modulated reflected autoregressive processes. In particular, he assumed that each transition of the Markov chain generates a new interarrival time $A_{n+1}$ and its corresponding service time $S_{n}$, thus, considered the Markov-dependent version of the process analysed in [12]. Note that the specific case of $a=1$ in [20] corresponds to the waiting time in a single server queue with Markov-dependent interarrival and service times studied in [1]. Markov-dependent structure of the form considered in [1] has also been used in insurance mathematics; see [4]. The process analysed in [1] (i.e., for $a=1$) is a special case of the class of processes studied in [6], although in [1], all the results were given explicitly. The case where a=1a=-1 was investigated in [31] (see also [32, Chapter 5]) in the context of carousel models. In [20] the author focused on the case where $a\in(0,1)$. Moreover, contrary to the case considered in [1, 31], in which given the state of the Markov chain at times $n$, $n+1$, the distributions of $A_{n+1}$, $S_{n}$ are independent of one another for all $n$ (although their distributions depend on the state of the background Markov chain), he further considered the case where there is a dependence among $\{S_{n}\}_{n\in\mathbb{N}_{0}}$, $\{A_{n}\}_{n\in\mathbb{N}}$ based on Farlie-Gumbel-Morgenstern (FGM) copula. More precisely, $\{(S_{n},A_{n+1})\}_{n\in\mathbb{N}_{0}}$ form a sequence of i.i.d. random vectors with a distribution function defined by FGM copula and dependent on the state of the underlying discrete time Markov chain. He also considered the case where $\{(S_{n},A_{n+1})\}_{n\in\mathbb{N}_{0}}$ have a bivariate matrix-exponential distribution, which is dependent on the state of the underlying discrete time Markov chain, as well as the case where there is a linear dependence among them. Special treatment was also given to the analysis of the case where the service times were dependent on the waiting time, as well as to the case where the server’s speed is workload proportional, i.e., a modulated shot-noise queue. The time-dependent analysis of a Markov-modulated reflected autoregressive process was also investigated.
@@ 1.2 Our contribution
Our primary goal is to explore a class of vector-valued stochastic recursions (6), in which some independence assumptions are lifted and for which, nevertheless, a detailed exact analysis can be provided. More precisely, by coping with the transient analysis of recursions (6), we result in a vector-valued functional equations of the form in (1), thus, extending considerably the work in [20], where vector-valued functional equations of the form in (5) were investigated, as well as the seminal works in [2, 12], where scalar functional equations of the form in (1) and (5) were studied, respectively. To solve this equation we make use of Liouville’s theorem [30] and Wiener-Hopf boundary value theory [15], although the vector-valued form requires additional technicalities. We also considered the stationary analysis of a simplified version of (6), that lead to a functional equation of type given in (2). We show that the solution of such type of equation does not necessarily requires the use of Liouville’s theorem and the reduction to a Wiener-Hopf boundary value problem. To cope with this kind of functional equations, we use as a vehicle Markov-dependent reflected autoregressive processes and other related models, for which the analysis results in functional equations of the form in (1), (2). We also go one step further and studied and consider extensions, under which given the state of the background Markov chain there is additional dependence structure based on the Farlie-Gumbel-Morgenstern (FGM) copula, i.e., $\tilde{Y}_{n}(X_{n})$, $\tilde{B}_{n}(X_{n})$ are not conditionally independent, but instead, they are dependent based on the FGM copula. In this work, we have mainly restricted ourselves to the case where $\alpha_{i}(s)$ are commutative contraction mappings. In particular, we focus on the case where $\alpha_{i}(s):=a_{i}s$. However, in section 6 we deal with a commutative mapping that is not a contraction, i.e., $\alpha_{i}(s)=s+\mu_{i}c$, $i=1,\ldots,N$.
We further consider a Markov-modulated fluid flow model with consumption, by generalising existed work in fluid queues by incorporating the concept of consumption. We also cope with a modulated ASIP tandem queue with Lévy input and consumption by extending the analysis of the model in [9] from the scalar to the Markov-modulated case, which in turn leads to a vector valued two-dimensional functional equation. Note that in such a case, we are dealing with the LST of the joint buffer content. In Section 8, we consider an even more general version of (2), related to the generating function of stationary queue-length distribution of a Markov-modulated $M/G/1$ queue with a general impatience scheme. This model is described by an integer-valued Markov-modulated reflected autoregressive process.
@@
Finally, in Section 9, we focus on the stationary analysis of a vector-valued autoregressive process (VAR(1)) that is described by recursions of the form $\tilde{Z}_{n+1}=[A\tilde{Z}_{n}+\tilde{S}_{n}-\tilde{A}_{n}]^{+}$, where now $A$ is an $N\times N$ known matrix. Our aim is to derive the LST of the stationary joint distribution of the process under study by solving a vector-valued functional equation of the form in (4). Despite the fact that the form of (4) is quite complicated, and the arguments of each component of the LST are linear combinations of $s_{i}$, $i=1,\ldots,N$, we succeed to solve it in case where the argument of $f(.)$ in the right hand side of (4) is an $N$-dimensional contraction mapping.

### v3

@@ Abstract
In this work, we study vector-valued functional equations with multiple recursive terms that arise naturally when we are dealing with vector-valued multiplicative Lindley-type recursions. We provide a detailed framework for the solution of such equations. Our theoretical results are applied in a wide range of semi-Markovian queueing, and vector-valued autoregressive processes.
Keywords: Vector-valued recursion; Laplace-Stieltjes transforms; Generating functions; Markov-modulation
@@ 1 Introduction
The primary aim of this work is to solve systems of functional equations of the form
[…]
for $Re(s)\geq 0$, where $\alpha_{i}(s)$, $i=1,\ldots,N$ are commutative contraction mappings. Equivalently, in matrix notation
where, $\tilde{Z}(s):=(z_{1}(s),\ldots,z_{N}(s))^{T}$, and
$H(s):=\begin{pmatrix}h_{1,1}(s)&h_{2,1}(s)&\ldots&h_{N,1}(s)\\ h_{1,2}(s)&h_{2,2}(s)&\ldots&h_{N,2}(s)\\ \ldots&\ldots&\ldots&\ldots\\ h_{1,N}(s)&h_{2,N}(s)&\ldots&h_{N,N}(s)\\ \end{pmatrix},\,\tilde{V}(s):=\begin{pmatrix}v_{1}(s)\\ v_{2}(s)\\ \vdots\\ v_{N}(s)\end{pmatrix}$,
are known matrix/vector functions, $\tilde{P}^{(i)}:=(\tilde{P}^{(i)})_{p,q}$, $i,p,q\in E=\{1,2,\ldots,N\}$ is an $N\times N$ matrix, with the element $\tilde{P}^{(i)}_{i,i}=1$, and all the other elements $\tilde{P}^{(i)}_{p,q}=0$, $p,q\neq i$. Note that $\sum_{i=1}^{N}\tilde{P}^{(i)}=I$, that is the identity matrix. Moreover, $\tilde{P}^{(i)}$ is a projection matrix onto the $i$th coordinate, $i\in E$. Such type of functional equations arise naturally when we are dealing with the stationary behaviour of vector-valued multiplicative Lindley-type recursions of a certain type with numerous applications in applied probability, including semi-Markovian queueing models, and vector-valued reflected autoregressive processes that are discussed in detail in the following sections. A queueing example of such a model is a Markov-modulated queue where arrivals and services are dependent on a background Markov chain, and the focus is on the workload just before an arrival that adds work in the system but makes obsolete a fixed portion of the work that is already in the system. This fixed portion also depends on the state of the background process and is related to the function $\alpha_{i}(s)$, $i\in E$.
To help the reader with the notation, unless otherwise specified, vectors/matrices are represented by uppercase letters, possibly augmented with diacritical marks such as a tilde or a hat, whereas their elements are denoted by lowercase letters.
The corresponding transient behaviour of such processes is described by similar, although a bit more complicated, functional equations of the form
@@
for $Re(s)\geq 0$, $Re(\eta)\geq 0$, $|r|<1$, and $\tilde{Z}(r,s,\eta):=(z_{1}(r,s,\eta),z_{2}(r,s,\eta),\ldots,z_{N}(r,s,\eta))^{T}$. Moreover,
$G(r,s,\eta)=r\begin{pmatrix}g_{1,1}(s,\eta)&g_{2,1}(s,\eta)&\ldots&g_{N,1}(s,\eta)\\ g_{1,2}(s,\eta)&g_{2,2}(s,\eta)&\ldots&g_{N,2}(s,\eta)\\ \ldots&\ldots&\ldots&\ldots\\ g_{1,N}(s,\eta)&g_{2,N}(s,\eta)&\ldots&g_{N,N}(s,\eta)\\ \end{pmatrix},\,\,K(r,s,\eta)=\begin{pmatrix}k_{1}(r,s,\eta)\\ k_{2}(r,s,\eta)\\ \vdots\\ k_{N}(r,s,\eta)\end{pmatrix}$.
We assume that $G(r,s,\eta):=rU(s,\eta)$, $K(r,s,\eta)$ are known matrix/vector functions.
A primary aim of this work is to extend the analysis in [2] to a matrix-valued framework. In particular, (1) corresponds to a vector-valued analogue of equation (2) in [2]. Functional equations of the form (1), (2) arise naturally from the analysis of multiplicative vector-valued stochastic recursions of the form
$\tilde{Z}_{n+1}=[V_{n}(Y_{n})\tilde{Z}_{n}+\tilde{S}_{n}(Y_{n})-\tilde{A}_{n}(Y_{n})]^{+},$ (3)
where $\{Y_{n};n\in\mathbb{N}\}$ denotes an irreducible Markov chain with a finite state space $E=\{1,2,\ldots,N\}$, and vectors $\tilde{Z}_{n},\tilde{S}_{n}(Y_{n}),\tilde{A}_{n}(Y_{n})\in\mathbb{R}^{n}$, and where $V_{n}(Y_{n}),\tilde{S}_{n}(Y_{n}),\tilde{A}_{n}(Y_{n})$ depend on the state of $Y_{n}$, i.e., we consider Markov-modulated vector-valued stochastic recursions.
Going one step further, our aim is also to deal with the solution of multidimensional versions of (1), which are quite challenging. A special case of such a class is given in (4),
@@
where $K(s,t)$, $R(s,t)$, $T(s)$, are known matrix-valued functions and $\alpha_{i}(t)$ are commutative contraction mappings in the right-half plane (e.g., $\alpha_{i}(t):=a_{i}t$, $a_{i}\in(0,1)$, $i=1,\ldots,N$). Note that this type of functional equation is different from (1) since now we have two coupled recursions: one in $t$ and one in $s$, which creates cross dependency between variables. However, the solution machinery is closely related to the one for (1). Vector-valued functional equations of the form (4) arise naturally in the modeling of Markov-modulated two-queue tandem queues with Lévy input and consumption (see [9] for the scalar case, where the contractions in the arguments on the right hand side do not depend on a background Markov chain). There, our focus is on deriving the Laplace-Stieltjes transform (LST) of the steady-state joint buffer content distribution, which is shown to satisfy a vector-valued functional equation of the form (4).
In view of multidimensional versions of functional equations (1), (2), it would be of great interest to consider vector-valued reflected autoregressive processes that are described by stochastic recursions of the form $\tilde{Z}_{n+1}=[A\tilde{Z}_{n}+\tilde{S}_{n}-\tilde{A}_{n}]^{+}$, where the autoregressive parameter $A$ is now an $N\times N$ known matrix with elements $a_{i,j}\in[0,1)$, and our main concern is to provide the joint LST of the stationary vector $\tilde{Z}$. Under specific assumptions, the vector of the marginal LSTs is shown to satisfy a (more general) functional equation of the form (1) (that can be handled in a similar way). On the other hand, in the case where our focus is on the joint LST of the stationary vector $\tilde{Z}$, the functional equation has a different form. In particular, when $N=2$ we have:
$\begin{array}[]{rl}f(s_{1},s_{2})=&c_{1}(s_{1},s_{2})f(T(s_{1},s_{2}))-c_{2}(s_{1},s_{2})f(T(s_{1},\lambda_{2}))\vskip 5.69054pt\\ &-c_{3}(s_{1},s_{2})f(T(\lambda_{1},s_{2}))+c_{4}(s_{1},s_{2})f(T(\lambda_{1},\lambda_{2})),\end{array}$ (5)
where, $T(s_{1},s_{2}):=A^{T}\tilde{s}=(a_{1,1}s_{1}+a_{2,1}s_{2},a_{1,2}s_{1}+a_{2,2}s_{2})$, $\tilde{s}=(s_{1},s_{2})^{T}$, and $c_{i}(s_{1},s_{2})$, $i=1,2,3,4$ are known functions. The form of (5) seems to be simpler compared to (1), since the autoregressive parameter (i.e., the matrix $A$) does not depend on the state of a background Markov chain, although we have a two-dimensional argument that is expressed as a linear combination of $s_{1}$, $s_{2}$ that complicates the analysis. However, in case $T(s_{1},s_{2})$ is a contraction mapping, its solution method follows steps similar to the one of (1), since it relies on the use of multiple recursive terms.
@@ 1.1 Related work
In the following, we provide a brief overview of the existing analytical results in the scalar case, since to the best of our knowledge, the only vector-valued version of the stochastic recursion (3) that has been treated analytically refers to the case where $V_{n}(Y_{n})=a\in(0,1)$; see [20].
@@
In [12], the authors considered the (scalar) recursion $Z_{n+1}=[aZ_{n}+S_{n}-A_{n+1}]^{+}$ (with $x]^{+}:=max\{x,0\}$), where $\{S_{n}-A_{n+1}\}_{n\in\mathbb{N}_{0}}$ forms a sequence of independent and identically distributed (i.i.d.) random variables and $a\in(0,1)$. The authors provided explicit results for the case where $\{S_{n}\}_{n\in\mathbb{N}_{0}}$ being a sequence of independent $exp(\lambda)$ distributed random variables, and $\{Y_{n}\}_{n\in\mathbb{N}_{0}}$ i.i.d., nonnegative and independent of $\{S_{n}\}_{n\in\mathbb{N}_{0}}$ with distribution function $F_{Y}(.)$ and Laplace–Stieltjes transform (LST) $\phi_{Y}(.)$. Note that in such a case, $Z_{n}$ could be interpreted as the workload in a queueing system just before the $n$th arrival, which adds $Y_{n}$ work, and makes obsolete a fixed fraction 1a1-a of the already present work. The case where $a=1$ corresponds to the classical Lindley recursion describing the waiting time of the classical M/G/1 queue, while the case where a=1a=-1 was investigated in [32]. Further progress has recently been made in [10] where the scalar autoregressive process described by the recursion $Z_{n+1}=[V_{n}Z_{n}+S_{n}-A_{n+1}]^{+}$ was investigated, where $\{V_{n}\}_{n\in\mathbb{N}_{0}}$ a sequence of i.i.d. random variables. In [11], the authors considered the case where $V_{n}Z_{n}$ replaced by $F(Z_{n})$, where $\{F(t)\}$ is a Levy subordinator (recovering also the case in [12], where $F(t)=at$). Recently, in [8], the authors motivated by applications that arise in queueing and insurance risk models, they considered Lindley-type recursions where the sequences $\{S_{n}\}_{n\in\mathbb{N}_{0}}$, $\{A_{n}\}_{n\in\mathbb{N}_{0}}$ obey a semi-linear dependence. These recursions can also be treated as of autoregressive type. Furthermore, the authors in [2] developed a method to study scalar versions of functional equations of the form (1) that arise in a wide range of queueing, autoregressive and branching processes. Finally, the author in [22], considered a generalized version of the model in [10], by assuming $V_{n}$ to take values in $(-\infty,1]$. In [3, eq. (9)], the authors considered a queueing system with two classes of impatient customers, leading to a specific example of (2) that was analyzed in detail; see also [24]. Quite recently, in [19] the authors generalized the work in [12] by considering, among others, non-trivial dependence structures among $\{S_{n}\}_{n\in\mathbb{N}_{0}}$, $\{A_{n}\}_{n\in\mathbb{N}}$. We also mention the work in [9] where, among others, the authors attempted to provide analytical results for ASIP (asymmetric inclusion process) tandem queues with consumption, the analysis of which leads to a functional equation that can be solved similarly as those in [11].
A primary motivation for our work is related to recent results on vector-valued reflected autoregressive processes. Quite recently, in [20], the author investigated vector-valued reflected autoregressive processes, where the sequences $\{S_{n}\}_{n\in\mathbb{N}_{0}}$, $\{A_{n}\}_{n\in\mathbb{N}}$ are governed by an irreducible background Markovian process with finite state space, i.e., the author considered Markov-modulated reflected autoregressive processes, where each transition of the Markov chain generates a new interarrival time $A_{n+1}$ and its corresponding service time $S_{n}$, thus, considered the Markov-dependent version of the process analysed in [12]. The analysis in [20] results in vector-valued functional equations of the form (6).
$\begin{array}[]{rl}\tilde{Z}(r,s,\eta)=&G(r,s,\eta)\tilde{Z}(r,as,\eta)+K(r,s,\eta),\vskip 5.69054pt\\ \tilde{Z}(s)=&H(s)\tilde{Z}(as)+\tilde{V}(s),\end{array}$ (6)
where $a\in(0,1)$.
@@
Note that the specific case of $a=1$ in [20] corresponds to the waiting time in a single server queue with Markov-dependent interarrival and service times studied in [1]. Markov-dependent structure of the form considered in [1] has also been used in insurance mathematics; see [4]. The process analysed in [1] (i.e., for $a=1$) is a special case of the class of processes studied in [6], although in [1], all the results were given explicitly. The case where a=1a=-1 was investigated in [31] (see also [32, Chapter 5]) in the context of carousel models. In [20] the author focused on the case where $a\in(0,1)$. Moreover, contrary to the case considered in [1, 31], in which given the state of the Markov chain at times $n$, $n+1$, the distributions of $A_{n+1}$, $S_{n}$ are independent of one another for all $n$ (although their distributions depend on the state of the background Markov chain), the author further considered the case where there is a dependence among $\{S_{n}\}_{n\in\mathbb{N}_{0}}$, $\{A_{n}\}_{n\in\mathbb{N}}$ based on Farlie-Gumbel-Morgenstern (FGM) copula. More precisely, $\{(S_{n},A_{n+1})\}_{n\in\mathbb{N}_{0}}$ form a sequence of i.i.d. random vectors with a distribution function defined by FGM copula and dependent on the state of the underlying discrete time Markov chain. The author also considered the case where $\{(S_{n},A_{n+1})\}_{n\in\mathbb{N}_{0}}$ have a bivariate matrix-exponential distribution, which is dependent on the state of the underlying discrete time Markov chain, as well as the case where there is a linear dependence among them. Special treatment was also given to the analysis of the case where the service times were dependent on the waiting time, as well as to the case where the server’s speed is workload proportional, i.e., a modulated shot-noise queue. The time-dependent analysis of a Markov-modulated reflected autoregressive process was also investigated.
@@ 1.2 Our contribution
Our primary goal is to explore a class of vector-valued functional equations (1), (2), (4) that arise naturally when we are dealing with Markov-modulated stochastic recursions. More precisely, vector-valued functional equations (1), (2) arise in the analysis of stochastic recursions (3). Equations of the form (4) arise when we are dealing with multi-dimensional versions of (3).
In Section 2, we provide a detailed theoretical framework in solving (1), (2) (see Theorems 1, 2), and thus extending the methodology developed in [2] from the scalar to the vector/matrix framework, also taking into account the transient behavior. We also provide a theoretical framework to solve (4); see Theorem 3. We emphasize that although (4) is of more general structure compared with (1), its solution machinery is similar.
The theoretical framework developed in Section 2 is applied in a number of application examples: More precisely, we first consider a MAP/G/1-type reflected autoregressive process, the analysis of which results in a functional equation of the type given in (1). We also go one step further and consider extensions, under which given the state of the background Markov chain, there is additional dependence structure based on the FGM copula, i.e., the service and the interarrival times are not conditionally independent, but instead they are dependent based on the FGM copula. In this work, we have mainly restricted ourselves to the case where $\alpha_{i}(s)$ are commutative contraction mappings. In most of the application examples, we focus on the case where $\alpha_{i}(s):=a_{i}s$. However, in subsection 3.3 we deal with a commutative mapping that is not a contraction as in [3], i.e., $\alpha_{i}(s)=s+\mu_{i}c$, $i=1,\ldots,N$. In subsection 3.4, we consider an even more general version of (1), related to the generating function of stationary queue-length distribution of a Markov-modulated $M/G/1$ queue with a general impatience scheme. This model is described by an integer-valued Markov-modulated reflected autoregressive process.
We then cope with the transient analysis of stochastic recursions (3) that results in a vector-valued functional equation of the form (2), thus, extending considerably the work in [20], where vector-valued functional equations of the form in (6) were investigated, as well as the seminal works in [2, 12], where scalar functional equations of the form (2), (1) were studied, respectively. More explicitly, we focus on a Markov modulated M/G/1-type reflected autoregressive process. The analysis of such a model leads to a vector-valued functional equation of the form (2). To come up with this equation, we make use of Liouville’s theorem [30] and Wiener-Hopf boundary value theory [15]. We further consider the transient analysis of a Markov-modulated fluid flow model with consumption by generalising existing work in fluid queues by incorporating the concept of consumption. That case leads also to a functional equation of the form (2).
Our next contribution relies on extending the analysis to multidimensional versions of (1), as given in (4). An interesting application example is a modulated ASIP tandem system of two queues with Lévy input and consumption. This is the Markov-modulated version of the model in [9], where we have further assumed that the consumption rate depends on the state of the background Markov chain. Thus, we considerably generalize [9]. Our analysis leads to a vector valued functional equation of the form (4) for the LST of the steady-state joint buffer content distribution.
@@
Finally, we investigate more intricate vector-valued functional equations than (1) by focusing on the stationary analysis of a vector-valued autoregressive process (VAR(1)) that is described by stochastic recursions of the form $\tilde{Z}_{n+1}=[A\tilde{Z}_{n}+\tilde{S}_{n}-\tilde{A}_{n}]^{+}$, where $\tilde{Z}_{n}=(Z_{1,n},\ldots,Z_{N,n})^{T}$, and $A$ is an $N\times N$ matrix. Under some independence assumptions, the vector of the marginal LSTs of the stationary version $\tilde{Z}$ satisfies a slightly more general functional equation of the form (1), although its solution method is similar. A quite challenging part is to consider the LST the stationary joint distribution of $\tilde{Z}$ that satisfies a vector-valued functional equation of the form (5). Despite the fact that the form of (5) is quite complicated, and the arguments of each component of the LST are linear combinations of $s_{i}$, $i=1,\ldots,N$, we succeed to solve it in case where the argument of $f(.)$ in the right hand side of (5) is an $N$-dimensional contraction mapping. Its solution method is similar to that of (1) and involves multiple recursive terms; see Theorem 32.